# Late-time transition of $M_{B}$ inferred via neural networks

### v1

@@ Abstract
The strengthening of tensions in the cosmological parameters has led to a reconsideration of fundamental aspects of standard cosmology. The tension in the Hubble constant can also be viewed as a tension between local and early Universe constraints on the absolute magnitude $M_{B}$ of Type Ia supernova. In this work, we reconsider the possibility of a variation of this parameter in a model-independent way. We employ neural networks to agnostically constrain the value of the absolute magnitude as well as assess the impact and statistical significance of a variation in $M_{B}$ with redshift from the Pantheon+ compilation, together with a thorough analysis of the neural network architecture. We find an indication for a transition redshift at the $z\approx 1$ region.
I Introduction
@@
The $\Lambda$CDM concordance model has provided a good description of both astrophysical and cosmological phenomenology for several decades Peebles:2002gy ; Copeland:2006wr . In this regime, cold dark matter (CDM) acts to stabilize galaxies Baudis:2016qwx ; XENON:2018voc while the late time accelerated expansion of the Universe Riess:1998cb ; Perlmutter:1998np is driven by a cosmological constant Mukhanov:991646 . The standard model of cosmology is completed by the inclusion of an inflationary epoch in the very early Universe Guth:1980zm ; Linde:1981mu . On the other hand, the concordance model has several underlying problems that includes the theoretical value of the cosmological constant Weinberg:1988cp , the UV completeness of the theory Addazi:2021xuf , and questions about the direct measurement of CDM particles LUX:2016ggv ; Gaitskell:2004gd . Recently, an observation-driven inconsistency has arisen in which the predictions of the $\Lambda$CDM model for the value of the Hubble constant (based on early Universe measurements) and the value obtained through surveys appear to diverge Abdalla:2022yfr . This Hubble tension is part of a larger cosmological tensions problem in which different cosmological parameters have differing predicted and observed values DiValentino:2020vhf ; DiValentino:2020zio ; DiValentino:2020vvd ; Staicova:2021ajb ; DiValentino:2021izs ; Perivolaropoulos:2021jda ; DiValentino:2022oon ; SajjadAthar:2021prg ; CANTATA:2021ktz ; Colgain:2022tql ; Krishnan:2021jmh ; Anderson:2023aga .
@@
There has been a dramatic increase in the reporting of the divergence of Hubble constant values in recent years. The discrepancy between direct and indirect measurements of $H_{0}$ has reached a critical threshold in statistical significance Aghanim:2018eyx ; ACT:2023kun ; Schoneberg:2022ggi . For the early Universe, the latest reported Hubble constant by the Planck and ACT collaborations are respectively $H_{0}^{\rm P18}=67.4\pm 0.5\,{\rm km\,s}^{-1}{\rm Mpc}^{-1}$ Aghanim:2018eyx and $H_{0}^{\rm ACT-DR4}=67.9\pm 1.5\,{\rm km\,s}^{-1}{\rm Mpc}^{-1}$ ACT:2020gnv . While in the late Universe, direct measurements have produced corresponding values by the SH0ES Team and the Carnegie Supernova Project gives respective reported values of $H_{0}^{\rm S22}=73.5\pm 1.1\,{\rm km\,s}^{-1}{\rm Mpc}^{-1}$ Brout:2022vxf and $H_{0}^{\rm CSP23}=73.22\pm 1.45\,{\rm km\,s}^{-1}{\rm Mpc}^{-1}$ Uddin:2023iob . Other reported values based on strong lensing by TDCOSMO gives a Hubble constant in the range $H_{0}^{\rm TDCOSMO23}=73.3\pm 5.8\,{\rm km\,s}^{-1}{\rm Mpc}^{-1}$ Shajib:2023uig . Each survey shows a high level of internal consistency, the tension arises between survey results, and is particularly poignant for surveys that rely on the $\Lambda$CDM concordance model in making their estimates Bernal:2016gxb .
@@
There has been a diverse range of reactions to the cosmological tensions problem with the prospect of systematics being exceedingly unlikely since the problem has appeared across a wide range of unconnected surveys. To tackle this growing issue, there have been a number of proposals of modifications to the concordance model in which the early Universe Poulin:2023lkg , the neutrino sector DiValentino:2021imh , or the underlying gravitational model Addazi:2021xuf ; CANTATA:2021ktz ; Cai:2019bdh ; LeviSaid:2021yat have been changed. Another approach that is gaining increased interest in the literature is that of using nonparametric approaches to reconstruct the evolution of cosmological parameters using methods that are independent of physical models. Currently, the most studied implementation of these approaches has taken the form of Gaussian processes 10.5555/1162254 which is based on the use of a kernel function that characterizes the covariance relationships between the observed data. The kernels are defined on nonphysical hyperparameters which can then be fit using traditional techniques. There have been a plethora of works in this direction Busti:2014aoa ; Busti:2014dua ; Seikel:2013fda ; Bernardo:2021mfs ; Yahya:2013xma ; 2012JCAP…06..036S ; Shafieloo:2012ht ; Benisty:2020kdt ; Benisty:2022psx ; Bernardo:2021cxi ; Bernardo:2022pyz ; Escamilla-Rivera:2021rbe ; Mukherjee:2021epjc ; Mukherjee:2020vkx ; Shah:2023rqb ; Mukherjee:2023lqr , which includes the reconstruction of various classes of cosmological models Cai:2019bdh ; Bernardo:2021qhu ; Ren:2022aeo ; Briffa:2020qli ; LeviSaid:2021yat . These approaches are interesting but as an approach, Gaussian processes possess several shortfalls in that the reconstructions it produce seem to have an overfitting issue at low values of redshift, as well as an over-reliance on the form of the kernel function.
@@
Another nonparameteric approach that has shown promise in confronting the problem of removing model reliance in data is that of artificial neural network (ANN) 10.2307/j.ctt4cgbdj ; Escamilla-Rivera:2019hqt where the number of nonphysical hyperparameters is dramatically increased and where there is no appearance of the problematic kernel covariance function. In this method, neurons are modeled on their biological analogs, which are then organized into layers that pass signals or inputs through the entirety of the network. This could directly take the form of redshift inputs and Hubble parameter outputs aggarwal2018neural ; Wang:2020sxl ; Gomez-Vargas:2021zyl . One example of an ANN implementation in the cosmological context is Ref. Wang:2019vxv where the Hubble diagram was reconstructed for a number of data sets, this was then further developed in Ref. Dialektopoulos:2021wde to include more complexity in the training data as well as null test analysis of the reconstructions. In Ref. Dialektopoulos:2023dhb this was further solidified with the native inclusion of the underlying covariances in cosmological data forming part of the training process in which the Hubble diagram is produced. ANNs have also been used to devise an approach in which the derivatives of reconstructed parameters, and their associated uncertainties, can be calculated Mukherjee:2022yyq which has led the way to making these methods competitive with Gaussian processes in that they can be used to reconstruct classes of modified cosmological models Dialektopoulos:2023jam .
@@
In this work, we use the robustness of reconstructions based on ANN architectures to probe the potential evolution of the absolute magnitude of type Ia supernovae events (SNIa). It has been suggested in the literature that the Hubble tension may be interpreted as a tension in the absolute magnitude Camarena:2022iae ; Camarena:2021jlr . Indeed, one can certainly transform the Hubble tension into a tension in this value; however, this may imply new physics at the level of the behaviour of astrophysical events. A consequence of this tension in the absolute magnitude would be that one cannot use priors in their analysis of modified cosmological models Efstathiou:2021ocp . Another consequence of this perspective of the tension in the expansion rate of the Universe at present is that one may devise proposals of models through which the absolute magnitude may vary in redshift space. Through this viewpoint, the instrument of nonparametric reconstruction techniques can be used to probe the possible evolution of the absolute magnitude in its own right. Model-independent constraints on the absolute magnitude was attempted by Dinda & Banerjee Dinda:2022jih employing Gaussian processes (GP), where an extensive analysis was carried out with the choice of different kernels and mean functions. As a rigorous extension of Dinda:2022jih , joint constraints on some cosmological nuisance parameters were attempted in Refs. Dinda:2022vmb ; Dinda:2023svr employing different combinations of data sets. Similar efforts in these directions were carried out in Ref. Banerjee:2023evd ; Mukherjee:2021kcu ; Mukherjee:2021epjc . Recent attempts to explore the inherent degeneracy between the parameters, $M_{B}$ and $r_{d}$ (comoving sound horizon at drag epoch) with deep neural networks, was carried out in Ref. Shah:2024slr . Indeed, there have been several physically motivated proposals for the potential evolution of the absolute magnitude parameter coming from phenomena. In Ref. Benisty:2022psx , the constancy of the absolute magnitude was probed comparatively through the reconstruction of the absolute magnitude using three general methods composed of ANNs, two kernel choices of GP reconstruction, and four models taken from the literature Tutusaus:2017ibk ; 2009A&A…506.1095L . Similar efforts to constraining $M_{B}$ or revisiting the constancy of $M_{B}$ in a non-flat universe were carried out by Mukherjee:2020vkx ; Favale:2023lnp for different combinations of datasets. However, results showed no significant evidence to merit an evolution in this parameter. It deserves mention that some of these non-parametric studies employed a combined approach in which baryon acoustic oscillation data (BAO) was used to isolate the luminosity distance and thus infer a value of the absolute magnitude. However, the BAO data is not completely model-independent and may have inserted a preference for this conclusion.
@@
In the current work, we are motivated to reconsider the uncalibrated latest Type Ia supernovae data, namely the Pantheon+ compilation, and use this to reconstruct the apparent magnitudes and their derivatives in a cosmology-independent way using neural networks. Through this route, we can constrain the value of the absolute magnitude $M_{B}$ without making any assumptions on the underlying cosmological model whatsoever. This is achieved by comparing the cosmic chronometer (CC) measurements of the Hubble parameter with the ANN reconstructed Hubble diagram from Pantheon+, thereby computing the absolute magnitude value in an agnostic way. We do this by introducing the technical details of the ANN approach in Sec. II, which directly leads to the reconstruction methodology adopted as detailed in Sec. III. In Sec. IV, we show the cosmology-independent constraints on $M_{B}$ that ANNs can achieve using this approach. Finally, we give a summary of our main conclusions and future work in Sec. V.
II Artificial Neural Networks
@@
In this section, we describe briefly the method by which ANN architectures 2015arXiv151107289C have been adopted. Inspired from biological neural networks, ANNs are built as a collection of neurons which are organized into layers. An input layer, is usually used to enter the data into the network, an output layer is used to extract the parameters we want to learn their behaviour, while there is a series of consecutive layers of neurons in between, usually called hidden layers, that depend on a number of hyperparameters and are being optimized to best mimic the real data processes.
@@
Neurons are triggered by an activation function that, across the larger number of neurons, can be used to model the complex relationships that feature in the data. In our work, we used the Exponential Linear Unit (ELU) 2015arXiv151107289C as the activation function, specified by
$f(x)=\begin{cases}{x}&\text{if }x>0\\ {\alpha(e^{x}-1)}&\text{if }x\leq 0\end{cases}\,,$ (1)
@@
where $\alpha$ is a positive hyperparameter that controls the value to which an ELU saturates for negative net inputs, which we set to unity. Besides being continuous and differentiable, this function does not act on positive inputs while negative inputs tend to be closer to unity for more negative input values. The choice of the activation function for the hidden layers usually has to do with the type of the network, i.e. multilayer perceptron, convolutional, recurrent, etc, while the choice for the output layer activation function has to do with the type of the problem (classification or regression); more details can be found in DeepLearning-book .
@@
The linear transformations and activation function produce a huge number of so-called hyperparameters which are nonphysical and can be optimized using training data so that the larger system mimics some physical process as closely as possible. Indeed, the full number of hyperparameters is even larger since there are additional hyperparameters related to the connection of the neurons which drastically increases the freedom of the ANN systems. The process by which training takes place involves these hyperparameters being optimized by comparing the predicted result $\hat{\Upsilon}$ with the real values $\Upsilon$ (training data) so that their difference is minimized. This minimization is called the loss function. This function is minimized by fitting methods, such as gradient descent, which fixes the hyperparameters for specific data sets, or a combination of them. In our work, we adopt Adam’s algorithm 2014arXiv1412.6980K as our optimizer, which is a modification of the gradient descent method that has been shown to accelerate convergence.
@@
The direct absolute difference between the predicted ($\hat{\Upsilon}$) and training ($\Upsilon$) outputs summed for every redshift is called the L1 loss function. This is akin to the log-likelihood function for uncorrelated data in a Markov Chain Monte Carlo (MCMC) analysis and is a very popular choice for ANN architectures. There are other choices such as the mean squared error (MSE) loss function which minimizes the square difference between $\hat{\Upsilon}$ and $\Upsilon$, and the smooth L1 (SL1) loss function which uses a squared term if the absolute error falls below unity and absolute term otherwise; one can see a comparison of different loss functions in Dialektopoulos:2021wde . Thus, it is at the level of the loss function that complexities in the data are inserted into the ANN through optimization in the training of the numerous hyperparameters that make up the system. In our work, we do this by assuming a loss function that is more akin to correlated data when considering log-likelihood functions for MCMC analyses. To that end, we incorporate the covariance matrix $\mathcal{C}$ of a data set by taking a loss function, given as
${\rm L_{\chi^{2}}}=\sum_{i,j}\left[m_{\rm obs}(z_{i})-m_{\rm pred}(z_{i})\right]^{\text{T}}\mathcal{C}_{ij}^{-1}\left[m_{\rm obs}(z_{j})-m_{\rm pred}(z_{j})\right]\,,$ (2)
@@
ANNs that feature at least one hidden layer can approximate any continuous function for a finite number of neurons, provided the activation function is continuous and differentiable HORNIK1990551 , which means that ANNs are applicable to the setting of cosmological data sets. In this work, we utilize the code for reconstructing functions from data called Reconstruct Functions with ANN (ReFANN [^1] https://github.com/Guo-Jian-Wang/refann) Wang:2019vxv which is based on PyTorch [^2] https://pytorch.org/docs/master/index.html. The code was run on GPUs to speed up the computational time, as well as making use of batch normalization 2015arXiv150203167I prior to every layer which further accelerates the convergence.
III Reconstruction Methdology

### v2

# A possible late-time transition of $M_{B}$ inferred via neural networks

@@ Abstract
The strengthening of tensions in the cosmological parameters has led to reconsidering the fundamental aspects of standard cosmology. The tension in the Hubble constant can also be viewed as a tension between local and early Universe constraints on the absolute magnitude $M_{B}$ of Type Ia supernova. In this work, we reconsider the possibility of a variation of this parameter in a model-independent way. We employ neural networks to agnostically constrain the value of the absolute magnitude as well as assess the impact and statistical significance of a variation in $M_{B}$ with redshift from the Pantheon+ compilation, together with a thorough analysis of the neural network architecture. We find an indication for a possible transition redshift at the $z\approx 1$ region.
1 Introduction
@@
The $\Lambda$CDM concordance model has provided a good description of both astrophysical and cosmological phenomenology for several decades [1, 2]. In this regime, cold dark matter (CDM) acts to stabilize galaxies [3, 4] while the late time accelerated expansion of the Universe [5, 6] is driven by a cosmological constant [7]. The standard model of cosmology is completed by the inclusion of an inflationary epoch in the very early Universe [8, 9]. On the other hand, the concordance model has several underlying problems that includes the theoretical value of the cosmological constant [10], the UV completeness of the theory [11], and questions about the direct measurement of CDM particles [12, 13]. Recently, an observation-driven inconsistency has arisen in which the predictions of the $\Lambda$CDM model for the value of the Hubble constant (based on early Universe measurements) and the value obtained through surveys appear to diverge [14]. This Hubble tension is part of a larger cosmological tensions problem in which different cosmological parameters have differing predicted and observed values [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
@@
There has been a dramatic increase in the reporting of the divergence of Hubble constant values in recent years. The discrepancy between direct and indirect measurements of $H_{0}$ has reached a critical threshold in statistical significance [27, 28, 29]. For the early Universe, the latest reported Hubble constant by the Planck and ACT collaborations are respectively $H_{0}^{\rm P18}=67.4\pm 0.5\,{\rm km\,s}^{-1}{\rm Mpc}^{-1}$ [27] and $H_{0}^{\rm ACT-DR4}=67.9\pm 1.5\,{\rm km\,s}^{-1}{\rm Mpc}^{-1}$ [30]. While in the late Universe, direct measurements have produced corresponding values by the SH0ES Team and the Carnegie Supernova Project gives respective reported values of $H_{0}^{\rm S22}=73.5\pm 1.1\,{\rm km\,s}^{-1}{\rm Mpc}^{-1}$ [31] and $H_{0}^{\rm CSP23}=73.22\pm 1.45\,{\rm km\,s}^{-1}{\rm Mpc}^{-1}$ [32]. Other reported values based on strong lensing by TDCOSMO gives a Hubble constant in the range $H_{0}^{\rm TDCOSMO23}=73.3\pm 5.8\,{\rm km\,s}^{-1}{\rm Mpc}^{-1}$ [33]. Each survey shows a high level of internal consistency, the tension arises between survey results, and is particularly poignant for surveys that rely on the $\Lambda$CDM concordance model in making their estimates [34].
There has been a diverse range of reactions to the cosmological tensions problem with the prospect of systematics being exceedingly unlikely since the problem has appeared across a wide range of unconnected surveys. To tackle this growing issue, there have been a number of proposals of modifications to the concordance model in which the early Universe [35], the neutrino sector [36], or the underlying gravitational model [37, 11, 23, 38] have been changed. Nonetheless, the absence of a universally accepted cosmological model makes linking physical theory to current observational data challenging. To address this, researchers construct viable cosmological models directly from observational data, reversing the traditional method of deriving evolution from Einstein’s field equations. This reverse-engineering approach aka reconstruction utilizes observational data to directly infer the distribution of matter sector, a technique widely used in modern cosmology. Key quantities often studied include the dark energy equation of state parameter $w_{d}$ [39, 40] and the quintessence potential $V(\phi)$ [41, 42, 43]. The direct approach is a parametric reconstruction where the cosmological quantities are represented as simple functions of redshift, along with some model parameters that are estimated using observational data [43, 44, 45, 46, 47, 48, 49]. However, this method can introduce bias due to its reliance on a specific functional form for the redshift dependence.
@@
Another approach that is gaining increased interest in the literature is that of using nonparametric approaches to reconstruct the evolution of cosmological parameters using methods that are independent of physical models. Currently, the most studied implementation of these approaches has taken the form of Gaussian processes [50] which is based on the use of a kernel function that characterizes the covariance relationships between the observed data. The kernels are defined on nonphysical hyperparameters which can then be fit using traditional techniques. There have been a plethora of works in this direction, which includes the reconstruction of various classes of cosmological models [51, 52, 53, 54, 55, 56, 57, 58, 37, 59, 60, 61, 62, 63, 38, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76]. These approaches are interesting, but as an algorithm, GPs possess several shortfalls such as overfitting issue at low redshifts, as well as an over-reliance on the form of the kernel function. As an alternative to GPs, Genetic Algorithms (GAs) have also been widely utilized for non-parametric reconstruction of cosmological functions using analytical forms as priors [77, 78, 79, 80, 81, 82]. Besides, methods like the iterative smoothing or local regression [83, 84, 85, 86, 87], weighted polynomial regression [88, 89], principal component analysis [90, 91, 92, 93], nodal and free-form reconstruction methods [94, 95, 96, 97, 98, 99, 100] have gained traction over the recent years. Although each of these methods has its limitations, particularly the interpretation of uncertainties can be more complex compared to GPs.
@@
Recently, machine learning approaches that have shown promise in confronting the problem of removing model reliance in cosmological data are artificial neural networks (ANN) [101, 102] where the number of nonphysical hyperparameters is dramatically increased. Besides, there is no appearance of the kernel covariance function compared to GPs. In this method, neurons are modeled on their biological analogs, which are then organized into layers that pass signals or inputs through the entirety of the network. This could directly take the form of redshift inputs and Hubble parameter outputs [103, 104, 105]. One example of an ANN implementation in the cosmological context is Ref. [106] where the Hubble diagram was reconstructed for a number of data sets, this was then further developed in Ref. [107] to include more complexity in the training data as well as null test analysis of the reconstructions. In Ref. [108] this was further solidified with the native inclusion of the underlying covariances in cosmological data forming part of the training process in which the Hubble diagram is produced. ANNs have also been used to devise an approach in which the derivatives of reconstructed parameters, and their associated uncertainties, can be calculated [109] which has led the way to making these methods competitive with Gaussian processes in that they can be used to reconstruct classes of modified cosmological models [110]. Recently, LADDER (Learning Algorithm for Deep Distance Estimation and Reconstruction) [111] has utilized deep neural networks (DNN) to reconstruct the cosmic distance ladder, highlighting the potential benefits of advanced learning techniques in enhancing optimal information extraction from cosmological data.
@@
In this work, we use the robustness of reconstructions based on ANN architectures to probe the potential evolution of the absolute magnitude of Type Ia supernovae events (SNIa). It has been suggested in the literature that the Hubble tension may be interpreted as a tension in the absolute magnitude [112, 113]. Indeed, one can certainly transform the Hubble tension into a tension in this value; however, this may imply new physics at the level of the behaviour of astrophysical events. A consequence of this tension in the absolute magnitude would be that one cannot use priors in their analysis of modified cosmological models [114]. Another consequence of this perspective of the tension in the expansion rate of the Universe at present is that one may devise proposals of models through which the absolute magnitude may vary in redshift space. Through this viewpoint, the instrument of nonparametric reconstruction techniques can be used to probe the possible evolution of the absolute magnitude in its own right. Model-independent constraints on the absolute magnitude was attempted by Dinda & Banerjee [115] employing Gaussian processes (GP), where an extensive analysis was carried out with the choice of different kernels and mean functions. As a rigorous extension of [115], joint constraints on some cosmological nuisance parameters were attempted in Refs. [116, 117] employing different combinations of data sets. Similar efforts in these directions were carried out in Ref. [61, 118, 119]. Recent attempts to explore the inherent degeneracy between the parameters, $M_{B}$ and $r_{d}$ (comoving sound horizon at drag epoch) with deep neural networks, was carried out in Ref. [111]. Indeed, there have been several physically motivated proposals for the potential evolution of the absolute magnitude parameter coming from phenomena. In Ref. [68], the constancy of the absolute magnitude was probed comparatively through the reconstruction of the absolute magnitude using three general methods composed of ANNs, two kernel choices of GP reconstruction, and four models taken from the literature [120, 121]. Similar efforts to constraining $M_{B}$ or revisiting the constancy of $M_{B}$ in a non-flat universe were carried out by [67, 122] for different combinations of datasets. However, results showed no significant evidence to merit an evolution in this parameter. It deserves mention that some of these non-parametric studies employed a combined approach in which baryon acoustic oscillation (BAO) data was used to isolate the luminosity distance and thus infer a value of the absolute magnitude. However, the BAO data is not completely model-independent and may have inserted a preference for this conclusion.
@@
In the current work, we are motivated to reconsider the uncalibrated latest Type Ia supernovae data, namely the Pantheon+ compilation, and use this to reconstruct the apparent magnitudes and their derivatives in a cosmology-independent way using neural networks. Through this route, we can constrain the value of the absolute magnitude $M_{B}$ without making any assumptions on the underlying cosmological model whatsoever. This is achieved by comparing the cosmic chronometer (CC) measurements of the Hubble parameter with the ANN reconstructed Hubble diagram from Pantheon+, thereby computing the absolute magnitude value in an agnostic way. We do this by introducing the technical details of the ANN approach in Sec. 2, which directly leads to the reconstruction methodology adopted as detailed in Sec. 3. In Sec. 4, we show the cosmology-independent constraints on $M_{B}$ that ANNs can achieve using this approach. Finally, we give a summary of our main conclusions and future work in Sec. 5.
2 Artificial Neural Networks
@@
This section briefly describes the method by which ANN architectures [123] have been adopted. Inspired from biological neural networks, ANNs are built as a collection of neurons that are organized into layers. An input layer is usually used to enter the data into the network, an output layer is used to extract the parameters we want to learn their behaviour, while there is a series of consecutive layers of neurons in between, usually called hidden layers, that depend on a number of hyperparameters and are being optimized to best mimic the real data processes.
@@
Neurons are triggered by an activation function that, across the larger number of neurons, can be used to model the complex relationships that feature in the data. In our work, we used the Exponential Linear Unit (ELU) [123] as the activation function, specified by
$f(x)=\begin{cases}{x}&\text{if }x>0\\ {\alpha(e^{x}-1)}&\text{if }x\leq 0\end{cases}\,,$ (2.1)
@@
where $\alpha$ is a positive hyperparameter that controls the value to which an ELU saturates for negative net inputs, which we set to unity. Besides being continuous and differentiable, this function does not act on positive inputs while negative inputs tend to be closer to unity for more negative input values. The choice of the activation function for the hidden layers usually has to do with the type of the network, i.e. multilayer perceptron, convolutional, recurrent, etc, while the choice for the output layer activation function has to do with the type of the problem (classification or regression); more details can be found in [124].
@@
The linear transformations and activation function produce a huge number of so-called hyperparameters which are nonphysical and can be optimized using training data so that the larger system mimics some physical process as closely as possible. Indeed, the full number of hyperparameters is even larger since there are additional hyperparameters related to the connection of the neurons which drastically increases the freedom of the ANN systems. The process by which training takes place involves these hyperparameters being optimized by comparing the predicted result $\hat{\Upsilon}$ with the real values $\Upsilon$ (training data) so that their difference is minimized. This minimization is called the loss function. This function is minimized by fitting methods, such as gradient descent, which fixes the hyperparameters for specific data sets, or a combination of them. In our work, we adopt Adam’s algorithm [125] as our optimizer, which is a modification of the gradient descent method that has been shown to accelerate convergence.
@@
The direct absolute difference between the predicted ($\hat{\Upsilon}$) and training ($\Upsilon$) outputs summed for every redshift is called the L1 loss function. This is akin to the log-likelihood function for uncorrelated data in a Markov Chain Monte Carlo (MCMC) analysis and is a very popular choice for ANN architectures. There are other choices such as the mean squared error (MSE) loss function which minimizes the square difference between $\hat{\Upsilon}$ and $\Upsilon$, and the smooth L1 (SL1) loss function which uses a squared term if the absolute error falls below unity and absolute term otherwise; one can see a comparison of different loss functions in [107]. Thus, it is at the level of the loss function that complexities in the data are inserted into the ANN through optimization in the training of the numerous hyperparameters that make up the system. In our work, we do this by assuming a loss function that is more akin to correlated data when considering log-likelihood functions for MCMC analyses. To that end, we incorporate the covariance matrix $\mathcal{C}$ of a data set by taking a loss function, given as
${\rm L_{\chi^{2}}}=\sum_{i,j}\left[m_{\rm obs}(z_{i})-m_{\rm pred}(z_{i})\right]^{\text{T}}\mathcal{C}_{ij}^{-1}\left[m_{\rm obs}(z_{j})-m_{\rm pred}(z_{j})\right]\,,$ (2.2)
@@
ANNs that feature at least one hidden layer can approximate any continuous function for a finite number of neurons, provided the activation function is continuous and differentiable [126], which means that ANNs are applicable to the setting of cosmological data sets. In this work, we utilize PyTorch [^1] https://pytorch.org/docs/master/index.html for our ANN implementation. The code was run on GPUs to speed up the computational time, as well as making use of batch normalization [127] prior to every layer which further accelerates the convergence.
3 Reconstruction Methodology